Shift report 20030119 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Jan 19 2003 - 16:16:16 EST

  • Next message: Apache: "Shift report 20030119 16:00-24:00"
    ------------------------------------------------------------
    Date: Sun Jan 19 16:16:16 2003, Shift: 08:00-16:00
    Supervisor: AM, with: 
    Log book#13 pages 93 - 94
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam was for the most part of the shift. Backgound noise
    is pretty low (Sc.15 ~ 10 kHz, Sc.16 ~ 20 kHz).
    2. Runs 7484, 7485, 7486, 7487 have been recorded.
    3. There is some problem with T3: in Supermon histogram specrtum "exists" just in first 8 wires... I sent an e-mail to Pawel with attached pics for T3, T4 amd T5 (for comparison) to ask for a solution to this.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sun Jan 19 2003 - 16:20:52 EST